OpenStudy (katherinesmith):

need help finding and graphing the focus and directrix of this equation

OpenStudy (katherinesmith):

the equation is x = -(1/12)y^2. i have already graphed it on paper but need to find the focus and directrix.

OpenStudy (tkhunny):

What's the vertex?

OpenStudy (katherinesmith):

i don't know how to find the vertex. that's my problem. after i find the vertex i know how to find the focus and directrix.

OpenStudy (tkhunny):

In the present form, the vertex is not too hard to find. Usually, \(x = -\dfrac{1}{4}y^{2}\) is decoded as \(x - 0 = -\dfrac{1}{4}(y-0)^{2}\) ans we see quickly that the Vertex is (0,0), The Origin.

OpenStudy (katherinesmith):

so what is p?

OpenStudy (tkhunny):

You need to find 4p from the rest of the information. In this case, -1/4.

OpenStudy (katherinesmith):

so the focus is (-1/4, 0) and the directrix is 1/4 ?

OpenStudy (tkhunny):

No, that's 1/(4p), not p. You need just p.

OpenStudy (katherinesmith):

so... 1/4(4) ? so 1?

OpenStudy (tkhunny):

Yes.

OpenStudy (katherinesmith):

so the focus is (-1,0) and the directrix is 1?

OpenStudy (tkhunny):

Since we ignored the sign, though, it's still up to you to decide which side of the vertex the focus will be found. Does this one open up, down, left, or right?

OpenStudy (katherinesmith):

it opens to the right so it's vertical so i'm pretty sure it's (-1, 0) and 1

OpenStudy (tkhunny):

I believe you have the focus and directrix correct, but it opens to the left or they are backwards. The focus has to be inside, right?

OpenStudy (tkhunny):

BTW "The directrix is 1" is not actually correct. The directrix is a line. It needs an equation of a line. The directrix is x = 1

OpenStudy (katherinesmith):

i know that. i was just summarizing.

OpenStudy (tkhunny):

Take the time to communicate clearly EVERY TIME. It will pay dividends now and great dividends later.

OpenStudy (katherinesmith):

@tkhunny can you help me with one more? \[y = \frac{ 1 }{ 4 }(x-2)^{2}\]

OpenStudy (anonymous):

no tu el borrito es P y minues 22 el conornoto